# Toward Green and Human-Like Artificial Intelligence: A Complete Survey on Contemporary Few-Shot Learning Approaches

### v2

@@ 4.4. Audio
In recent years, FSL has been successfully applied in the field of acoustic signal processing, with most applications focusing on audio classification(Wang and Anderson, 2022) and its extensions in the continual(Wang et al., 2021) and class-incremental(Li et al., 2023a) learning settings. Additionally, FSL has also been utilized for sound event detection(Wang et al., 2020a), while a promising application is that of speaker(Li et al., 2023a) and sound recognition (Chou et al., 2019), which can be effectively used in mobile devices and can benefit people with hearing difficulties.
4.5. Generative AI
@@
One of the most prominent and broadly expanding fields where FSL can be extremely useful is that of Generative AI. The ability to generate novel content given only a few reference samples has been extensively explored in the context of computer vision with recent applications, including image generation (Zheng et al., 2023a), style transfer(Tang et al., 2023a), and semantic image synthesis (Careil et al., 2023). FSL has also been utilized for generative purposes in different modalities such as time series(Jiang et al., 2022), video(Li et al., 2023b), speech(Kumar et al., 2020), and music(Liang et al., 2020) generation. Few-shot generation across different modalities has also gained much popularity recently via applications such as text-to-image(Zhou et al., 2022) and text-to-speech(Huang et al., 2022) generation.
4.6. Automated Machine Learning
@@
Traditionally, Automated Machine Learning (AutoML) has been one of the most used testbeds for novel meta-learning techniques. AutoML aims to automatically find the best model architecture and hyperparameters for a given problem and dataset. This reduces the need for domain-specific knowledge and extensive hyperparameter tuning, making machine learning more accessible and accelerating deployment. As for FSL in the context of AutoML, applications have mainly focused on Neural Architecture Search (Lee et al., 2023a) (NAS), model selection(Park et al., 2022), and hyperparameter optimization(Maraval et al., 2024).
4.7. Healthcare
@@
A natural field of application for FSL is when data collection can be difficult, costly, or even impossible. Consequently, healthcare is a perfect application field for FSL since most patient data is subject to privacy restrictions, yet high performance needs to be maintained under these critical settings. Recently, FSL has been successfully applied in medical image classification (Sun et al., 2023), medical image segmentation (Huang et al., 2023), and 3D medical image segmentation in a federated setting (He et al., 2023). FSL has also been critical in cases where more data acquisition is impossible, such as in rare disease classification (Sun et al., 2023), and drug discovery (Chen et al., 2022). Finally, some novel approaches have focused on developing models that can handle multiple downstream medical tasks such as operational outcomes, anticipating lab test values, assignment of new diagnoses, and anticipating chest X-ray findings(Wornow et al., 2023).
4.8. Recommendation
@@
A field that has particularly benefited from the use of FSL techniques is that of recommendation systems, by helping overcome major challenges, such as the cold-start problem, and improve model performance(Yu et al., 2021). Additionally, personalization is another area that has greatly benefited from FSL(Shysheya et al., 2023). Specifically, by treating each user’s data as a separate task, FSL can efficiently learn across all of them, even with limited data. Consequently, FSL has successfully been applied for personalized video retrieval(Yeh et al., 2023), personalized recommendation(Wu et al., 2022a), and personalization in federated learning settings(Zhao et al., 2022).
5. Recent Trends in Few-Shot Learning
Cross-Domain FSL
Cross-Domain FSL (CD-FSL), a learning paradigm combining FSL’s knowledge transfer to tasks with novel classes and Domain Adaptation’s knowledge transfer across different data distributions, has emerged in recent years. CD-FSL’s goal is to leverage knowledge extracted from base tasks to adapt to novel tasks with different classes that follow different distributions. A common challenge in CD-FSL is that representations on the target dataset are often biased due to the data volume discrepancy, making it difficult to obtain robust models across different domains. Although more challenging, this setting is particularly interesting since, in practice, it is common for the base and target datasets of FSL to originate from different domains, necessitating cross-domain evaluation. To address this issue, recent approaches have focused on mitigating cross-domain bias during both the development and evaluation stages.
At the development level, CD-FSL includes models that are explicitly trained to handle few-shot domain adaptation/generalization, with methods such as finetuning a model’s projection and prediction heads(Jang et al., 2023), retraining the feature extractor using an inverted Firth Bias regularizer(Wang et al., 2022b), and using knowledge distillation(Zheng et al., 2023b). The evaluation level includes few-shot cross-domain adaptation of models, even if they were not explicitly trained for this task, which is useful in a wide variety of applications such as visual question answering, temporal point processes, neural architecture search, and drug discovery(Schimunek et al., 2023).
Overall, robustness against domain shifts can be a potential step towards Broad AI(Schimunek et al., 2023). At the same time, cross-domain evaluation in meta-learning could help measure overfitting at the meta-level by attributing generalization to learning or memorization at the task level.
In-context Learning in Large Language Models
Although not explicitly trained for it, it has been demonstrated that LLMs can generalize to novel tasks when provided with just a few demonstrations(Brown et al., 2020). This few-shot ability is commonly referred to as in-context learning(Min et al., 2023) (ICL) since the few examples serve as a demonstration context that enables generalization. Although there have been some attempts to shed light on the mechanisms that enable ICL in LLMs, it is evident that there are some significant discrepancies between ICL and FSL. One of the major differences is that ICL is based on prompts that provide the context for the task through input-output pairs of examples, obviating the need for parameter finetuning via gradient backpropagation, thus significantly reducing computing power and time associated with novel task adaptation and enabling faster and more efficient training. Finally, while FSL models are explicitly trained for rapid adaptation to new tasks, ICL emerges naturally in LLMs and is also interpretable as demonstrations are in natural language. These desirable characteristics of ICL have undoubtedly rendered LLMs particularly attractive in various applications where data is limited(Dai et al., 2022). However, ICL also faces limitations related to its instability caused by different factors, such as prompting format, example content, and ordering. To mitigate these issues, different calibration techniques have been proposed, such as using prototypical networks to calibrate the decision boundary of LLMs(Han et al., 2023).
Finally, ICL has been combined with other FSL paradigms, notably meta-learning, to enhance its performance. By viewing ICL as a type of prompt learning, meta-learning can be used to efficiently learn better prompts across various downstream tasks. In this direction, recent approaches include meta-training on few-shot formulated tasks(Min et al., 2022), using MAML to learn a good prompt initialization(Hou et al., 2022), and meta-learning different priors and learning strategies across sets of downstream tasks(Coda-Forno et al., 2023).
Few-Shot Evaluation of Foundation Models
@@ 5. Recent Trends in Few-Shot Learning
Based on these observations and the recent bloom of foundation models, one recent trend has been the evaluation and adaptation of these models regarding their FSL capabilities. Few-shot and even zero-shot performance is increasingly assessed both in the language(Brown et al., 2020) and visual domain(Dosovitskiy et al., 2020), as well as in multimodal settings(Radford et al., 2021). Additionally, variants of foundation models not explicitly developed for FSL(Kwak et al., 2023), combinations of foundation models aiming to enhance few-shot performance(Zhang et al., 2023a) have also emerged.
@@
Few-Shot Federated Learning
In recent years, decentralized types of learning like Federated Learning (FL) have been researched extensively to leverage scattered data across machines while preserving privacy. FL reduces communication costs and enhances data privacy by sharing only model parameters between servers and clients, but its performance suffers when data is scarce. To enable robust application of FL under the FSL setting, several approaches combining FL with meta-learning have been proposed, such as adjusting popular meta-learning algorithms for FL(Chen et al., 2018), recasting FedAvg as a meta-learning algorithm for personalized FL(Jiang et al., 2019) and separately executing meta-learning’s optimization loops on the server and clients(Fallah et al., 2020).
In the last few years, there has been a plethora of different few-shot federated learning (FSFL) methods developed for applications, such as facial expression recognition(Shome and Kar, 2021), load forecasting(Tang et al., 2023b), 3D MRI medical image classification(He et al., 2023), fault diagnosis in industrial settings(Chen et al., 2023c), and NLP(Dong et al., 2022). FSFL is ideal for these types of problems since the data is typically scarce, heterogeneous, and dispersed in different machines, while it also satisfies the need for privacy preservation because of the sensitive nature of the involved data.
Continual FSL
Following the recent trend of bridging the gap between FSL and tangential fields of study, FSL has also been applied in the context of Continual Learning. Continual Learning refers to the ability of a model to learn from an infinite stream of data, acquiring new knowledge along the way while also preserving the one already gained. In the case of Continual FSL, each session introduces novel few-shot tasks, increasing the danger of overfitting to the new limited training data. Some notable attempts in addressing Continual FSL include extending MAML in the online learning setting(Finn et al., 2019), introducing a mixture of across-tasks shared parameters through a Dirichlet process mixture model to create a meta-learning model capable of incrementally learning heterogeneous tasks(Jerfel et al., 2019), and using Bayesian meta-learning with different meta-knowledge components encapsulating across-tasks knowledge to enable continual learning of sequentially available heterogeneous tasks(Wu et al., 2023a).
Additionally, much of the recent research on Continual FSL has focused on Class-Incremental FSL (CIFSL) (Zhang et al., 2023b) where new classes are introduced in each session, and the learning model should be able to infer from all classes seen up to that point. In (Song et al., 2023b), to facilitate generalization to novel classes on the incremental sessions, a contrastive supervised loss is used to ensure adequate base class separation. Knowledge distillation, a commonly used method in incremental learning, has also been applied to this setting by leveraging two teacher networks to enable adaptation to novel session classes(Zhao et al., 2023). Finally, various recently introduced methods for CIFSL draw inspiration from techniques and insights from other areas of machine learning, such as recursive analytical learning(Zhuang et al., 2023), neural collapse(Yang et al., 2023), and the Lottery Ticket Hypothesis(Kang et al., 2023b).
Meta-Learning in novel settings
@@
Traditionally, most FSL applications have revolved around problems related to computer vision and RL. While FSL has also been applied to different domains, such as NLP, and learning problems, such as unsupervised learning, these constitute well-studied areas within the machine learning community. However, there has been a trend in applying meta-learning in novel and less-explored setups in recent years. It is worth noting that while data efficiency is a central consideration in FSL, time efficiency in terms of fast adaptation to the existing data within each task is critical in many of these settings, elucidating why meta-learning is broadly adopted in these novel scenarios. For instance, meta-learning has been applied in various games such as zero-sum games, potential games, general sum multi-player games, and Stackelberg security games, showing significant improvements(Harris et al., 2022). Meta-learning an optimal initialization has also been applied in combinatorial optimization(Wang and Li, 2023) by recasting unsupervised learning for combinatorial optimization as a meta-learning problem, as well as few-shot knowledge graph completion(Wu et al., 2023c). Additionally, meta-learning has been extended to learning cognitive tasks(Miconi, 2023) by introducing another evolutionary level of learning that enables modification of the network in a self-referential manner. Finally, meta-learning has been used to efficiently discover optimizing functions without access to gradients(Lange et al., 2023) and to solve optimal transport problems(Amos et al., 2023) rapidly.
6. Challenges and Future Directions
Green AI via FSL
@@ 6. Challenges and Future Directions
A central characteristic of modern LLMs and foundation models is their scaling in the regime of billions of parameters. While this has led to breakthrough performance results, it has also led to a commensurate increase in their energy consumption and carbon emissions. The environmental damage associated with training and operating such models has been highlighted in 2023’s AI index report(Maslej et al., 2023), showing that their carbon emissions exceed those of an average person’s lifetime. Other studies (Patterson et al., 2021) have also focused on measuring the carbon footprint of LLMs, yet more needs to be done to mitigate their impact.
@@
A first step toward this direction has been the development of novel Green AI metrics that quantify these effects. For instance, in (Wu et al., 2022b), the carbon footprint of AI models is estimated as:
(5) $CO_{2}^{embodied}=\sum_{i}\frac{Time}{Lifetime}CO_{2}^{embodied}(AI_{system})(i)$
@@
where $Time$ refers to the time the model is utilized over the specified $Lifetime$. Other tools that enable measurement of model training environmental footprint have also been proposed. On the other hand, in (Henderson et al., 2020), a formula for calculating the total energy consumption is derived:
(6) $e_{total}=PUE\sum_{p}(p_{dram}e_{dram}+p_{cpu}e_{cpu}+p_{gpu}e_{gpu})$
@@
Approaches such as FL have been explored in the context of Green AI, yet they face limitations due to hardware constraints, scarcity of renewable energy sources for edge devices, and the need for communication among devices. However, when combined with FSL, they can reduce energy consumption on each device up to 41.2%(Cai et al., 2023). Overall, FSL can alleviate the need for extensive model selection and hyperparameter tuning and reduce the volume of available data needed to achieve strong generalization performance. However, more research is required to establish FSL as a feasible solution for Green AI.
Meta-Learning vs Transfer Learning debate
@@
While meta-learning has been the leading paradigm for FSL for many years, various recent results have cast doubt on its dominance. Specifically, feature reuse has been identified as the main factor for fast adaptation in MAML(Raghu et al., 2020) while recent findings show that finetuning with modern pretrained backbones can also form strong baselines(El Baz et al., 2021). Consequently, transfer learning has emerged as a capable competitor of meta-learning, with numerous simple baselines based on finetuning being proposed and demonstrating comparable performance against complex meta-learning architectures.
@@
A shared insight in this line of work is that learning strong, transferable representations is crucial for FSL(Wang et al., 2022b), and the quality of these representations can be affected by various factors such as model size(Tian et al., 2020), invariance and equivariance(Rizve et al., 2021), support set size(Luo et al., 2023), and number of pretraining classes(Dhillon et al., 2020). Overall, further research is required to better understand the importance of expressive and generalizable representations in FSL. An interesting direction would also be to examine FSL’s interplay with other learning paradigms with similar objectives, such as self-supervised learning(Ni et al., 2021), and incorporate techniques from these fields to obtain more generalizable representations(Tian et al., 2020).
@@
Moreover, the lack of a unified evaluation protocol has constituted a major issue in comparing meta-learning and finetuning models. It is argued that progress using meta-learning techniques might be illusory, with incongruities in the implementation details of various FSL methods(Dhillon et al., 2020), and lack of evaluation under domain shift(Chen et al., 2019), identified as two main challenges obstructing meaningful comparative evaluations. To mitigate these issues, a hardness metric has been proposed(Dhillon et al., 2020), as well as the use of novel datasets(Dumoulin et al., 2021). However, more research is needed to determine more datasets and a unified evaluation protocol for FSL methods.
Adoption of FSL in new modalities
@@
Traditionally, most FSL and meta-learning techniques have commonly been applied in the image domain(Snell et al., 2017) and for RL(Wang et al., 2016). However, since FSL aims to alleviate the adaptation problem in cases where data is scarce, its application could be fruitful in various other domains and data modalities.
@@
A common type of data underexplored in the FSL setting is tabular data. Only recently, some novel approaches have emerged, including using Conditional Batch Normalization to create novel tasks with tabular data(Sun et al., 2023) and applying prototypical networks on tabular data tasks created via unsupervised meta-learning(Nam et al., 2023). Time series data have also received meager attention until recently. Some notable first attempts in applying FSL in that modality include a meta-learning approach for few-shot forecasting in high-dimensional time series based on a single latent dynamic function conditioned in each task(Jiang et al., 2022), load forecasting via MAML-based models(Xu et al., 2024), and meta-learning neural processes with application on point processes(Bae et al., 2023).
@@
One of the main reasons hindering FSL progress outside the visual domain and RL is the limited number of FSL datasets for these domains and modalities. Although some novel FSL datasets, such as FS-MOL(Stanley et al., 2021) for few-shot drug discovery, have emerged, the lack of similar datasets in modalities such as time series necessitates developing and establishing standardized FSL benchmarks in these domains.
Fully generalizable FSL

### v3

@@ 5.4 Audio
In recent years, FSL has been successfully applied in the field of acoustic signal processing, with most applications focusing on audio classification [216] and its extensions in the continual [220] and class-incremental [126] learning settings. Additionally, FSL has also been utilized for sound event detection [218], while a promising application is that of speaker [126] and sound recognition [30], which can be effectively used in mobile devices and can benefit people with hearing difficulties.
5.5 Generative AI
@@
One of the most prominent and broadly expanding fields where FSL can be extremely useful is that of Generative AI. The ability to generate novel content given only a few reference samples has been extensively explored in the context of computer vision with recent applications, including image generation [255], style transfer [196], and semantic image synthesis [21]. FSL has also been utilized for generative purposes in different modalities such as time series [94], video [124], speech [113], and music [128] generation. Few-shot generation across different modalities has also gained much popularity recently via applications such as text-to-image [257] and text-to-speech [85] generation.
5.6 Automated Machine Learning
@@
Traditionally, Automated Machine Learning (AutoML) has been one of the most used testbeds for novel meta-learning techniques. AutoML aims to automatically find the best model architecture and hyperparameters for a given problem and dataset [12]. This reduces the need for domain-specific knowledge and extensive hyperparameter tuning, making machine learning more accessible and accelerating deployment. As for FSL in the context of AutoML, applications have mainly focused on Neural Architecture Search [120] (NAS), model selection [160], and hyperparameter optimization [138].
5.7 Healthcare
@@
A natural field of application for FSL is when data collection can be difficult, costly, or even impossible. Consequently, healthcare is a perfect application field for FSL since most patient data is subject to privacy restrictions, yet high performance needs to be maintained under these critical settings. Recently, FSL has been successfully applied in medical image classification [192], medical image segmentation [84], and 3D medical image segmentation in a federated setting [76]. FSL has also been critical in cases where more data acquisition is impossible, such as in rare disease classification [192], and drug discovery [27]. Finally, some novel approaches have focused on developing models that can handle multiple downstream medical tasks such as operational outcomes, anticipating lab test values, assignment of new diagnoses, and anticipating chest X-ray findings [224].
5.8 Recommendation
@@
A field that has particularly benefited from the use of FSL techniques is that of recommendation systems, by helping overcome major challenges, such as the cold-start problem, and improve model performance [243]. Additionally, personalization is another area that has greatly benefited from FSL [185]. Specifically, by treating each user’s data as a separate task, FSL can efficiently learn across all of them, even with limited data. Consequently, FSL has successfully been applied for personalized video retrieval [240], personalized recommendation [225], and personalization in federated learning settings [253].
6 Recent Trends and Future Directions in Few-Shot Learning
The rapid growth of FSL methodologies and applications has established it as a distinct area within the machine learning community. However, FSL is not an isolated field; various methodologies developed within FSL overlap with different areas of machine learning research. Consequently, FSL has been influenced by advancements and new challenges from both its own field and related research areas. For instance, the emergence of large foundation models has led to their application and evaluation in settings with limited available data. Additionally, more general research directions, such as developing models that exhibit traits of human-like intelligence and advancing AI to be more environmentally friendly, have also penetrated the FSL field in recent years. In this section, we provide a discussion of various trends that have been rising in recent years in the FSL field, as well as the existing and novel challenges that may shape future research directions in FSL.
6.1 Few-Shot Evaluation of Foundation Models
@@
Based on these observations and the recent bloom of foundation models, one recent trend has been the evaluation and adaptation of these models regarding their FSL capabilities [132]. Few-shot and even zero-shot performance is increasingly assessed both in the language [19] and visual domain [43], as well as in multimodal settings [167]. Additionally, variants of foundation models not explicitly developed for FSL [114] and combinations of foundation models aiming to enhance few-shot performance [251] have also emerged.
@@
6.2 Meta-Learning in novel settings
@@
Traditionally, most FSL applications have revolved around problems related to computer vision and RL. While FSL has also been applied to different domains, such as NLP, as seen in Section 5, and under different problem formulations, such as unsupervised learning, as seen in Section 4, these constitute well-studied areas within the machine learning community. However, there has been a trend in applying meta-learning in novel and less-explored setups in recent years. It is worth noting that while data efficiency is a central consideration in FSL, time efficiency in terms of fast adaptation to the existing data within each task is critical in many of these settings, elucidating why meta-learning is broadly adopted in these novel scenarios.
Meta-learning has been applied in various games such as zero-sum games, potential games, general sum multi-player games, and Stackelberg security games, showing significant improvements [74]. Meta-learning an optimal initialization has also been applied in combinatorial optimization [211] by recasting unsupervised learning for combinatorial optimization as a meta-learning problem, as well as few-shot knowledge graph completion [228]. Additionally, meta-learning has been extended to learning cognitive tasks [142] by introducing another evolutionary level of learning that enables modification of the network in a self-referential manner. Finally, meta-learning has been used to discover optimizing functions without access to gradients efficiently [117] and to solve optimal transport problems [5] rapidly.
6.3 Green AI via FSL
@@
A central characteristic of modern LLMs and foundation models is their scaling in the regime of billions of parameters. While this has led to breakthrough performance results, it has also led to a commensurate increase in their energy consumption and carbon emissions. The environmental damage associated with training and operating such models has been highlighted in 2023’s AI index report [139], showing that their carbon emissions exceed those of an average person’s lifetime. Other studies [164] have also focused on measuring the carbon footprint of LLMs, yet more needs to be done to mitigate their impact.
@@
A first step toward this direction has been the development of novel Green AI metrics that quantify these effects. For instance, in [227], the carbon footprint of AI models is estimated as:
$CO_{2}^{embodied}=\sum_{i}\frac{Time}{Lifetime}CO_{2}^{embodied}(AI_{system})(i)$ (5)
@@
where $Time$ refers to the time the model is utilized over the specified $Lifetime$. Other tools that enable measurement of model training environmental footprint have also been proposed. On the other hand, in [77], a formula for calculating the total energy consumption is derived:
$e_{total}=PUE\sum_{p}(p_{dram}e_{dram}+p_{cpu}e_{cpu}+p_{gpu}e_{gpu})$ (6)
@@
Approaches such as FL have been explored in the context of Green AI, yet they face limitations due to hardware constraints, scarcity of renewable energy sources for edge devices, and the need for communication among devices. However, when combined with FSL, they can reduce energy consumption on each device up to 41.2% [20]. Overall, FSL can alleviate the need for extensive model selection and hyperparameter tuning and reduce the volume of available data needed to achieve strong generalization performance. However, more research is required to establish FSL as a feasible solution for Green AI.
6.4 Meta-Learning vs Transfer Learning debate
@@
While meta-learning has been the leading paradigm for FSL for many years, various recent results have cast doubt on its dominance. Specifically, feature reuse has been identified as the main factor for fast adaptation in MAML [168] while recent findings show that finetuning with modern pretrained backbones can also form strong baselines [48]. Consequently, transfer learning has emerged as a capable competitor of meta-learning, with numerous simple baselines based on finetuning being proposed and demonstrating comparable performance against complex meta-learning architectures.
@@
A shared insight in this line of work is that learning strong, transferable representations is crucial for FSL [209], and the quality of these representations can be affected by various factors such as model size [199], invariance and equivariance [175], support set size [136], and number of pretraining classes [40]. Overall, further research is required to understand more clearly the importance of expressive and generalizable representations in FSL. An interesting direction would also be to examine FSL’s interplay with other learning paradigms with similar objectives, such as self-supervised learning [153] and incorporate techniques from these fields to obtain more generalizable representations [199].
@@
Moreover, the lack of a unified evaluation protocol has constituted a major issue in comparing meta-learning and finetuning models. It is argued that progress using meta-learning techniques might be illusory, with incongruities in the implementation details of various FSL methods [40], and lack of evaluation under domain shift [28], identified as two main challenges obstructing meaningful comparative assessments. To mitigate these issues, a hardness metric has been proposed [40], as well as the use of novel datasets [45]. However, more research is needed to determine more datasets and a unified evaluation protocol for FSL methods.
6.5 Impact of underlying distributions
One of the main differences between FSL and standard machine learning problems is the limited number of support data within each task, which complicates the precise estimation of the underlying data distribution. While meta-learning approaches compensate for it by extracting shared knowledge across different tasks, it is still unclear how to effectively achieve accurate estimates of the underlying task-generating process. Since FSL typically relies on utilizing a set of tasks, this setting introduces another distribution at the meta-level, specifically the distribution of tasks. Consequently, similar to how data augmentation has been extensively utilized in standard machine learning to obtain a better evaluation of the underlying data-generating distribution, similar techniques can also be utilized at the meta-level to increase across-task generalization by learning a more accurate representation of the underlying task-generating distribution. Towards that direction, in recent years, there has been substantial progress in methods that employ interpolation and augmentation techniques at the meta-level.
In the case of task augmentation, various techniques, including Conditional Batch Normalization [192], mixing support and query sets with Manifold Mixup to create new query sets [236], and adversarial training [208] have been employed to generate new tasks. On the other hand, task augmentation techniques typically include leveraging information from two existing tasks to create a new one. Some of the most recent approaches include using memory banks [83], interval bounds [36], Manifold Mixup [237], and expressive neural set functions [121] to encode and interpolate between tasks.
Yet, despite these practical implementations, we still lack a more enhanced comprehension of the influence of the underlying data distributions on FSL at a theoretical level. Specifically, only a single work [260] that focuses on the theoretical examination of FSL emphasizes the importance of careful modeling of both the task distribution and the data distribution within each task. Consequently, further research is needed to understand more clearly the connection between task and data distributions and how they can efficiently be leveraged in the context of FSL.
6.6 Adoption of FSL in new modalities
@@
Traditionally, most FSL and meta-learning techniques have commonly been applied in the visual domain [187] and for RL [213]. However, since FSL aims to alleviate the adaptation problem in cases where data is scarce, its application could be fruitful in various other domains and data modalities.
@@
A common type of data underexplored in the FSL setting is tabular data. Only recently, some novel approaches have emerged, including using Conditional Batch Normalization to create novel tasks with tabular data [192] and applying prototypical networks on tabular data tasks created via unsupervised meta-learning [150]. On a broader setting focusing on small tabular tasks, in [78], the authors have pretrained a transformer-based model on synthetically generated tasks using Bayesian Neural Networks and Structural Causal Models and showed that the model can adapt to downstream tabular tasks without any additional finetuning, similar to the way LLMs perform ICL. Time series data have also received insufficient attention until recently. Some notable first attempts in applying FSL in that modality include a meta-learning approach for few-shot forecasting in high-dimensional time series based on a single latent dynamic function conditioned in each task [94], load forecasting via MAML-based models [232], and meta-learning neural processes with application on point processes [11].
@@
Overall, one of the main reasons hindering FSL progress outside the visual domain and RL is the limited number of FSL datasets for these domains and modalities. Although some novel FSL datasets, such as FS-MOL [190] for few-shot drug discovery, have emerged, the lack of similar datasets in modalities such as time series necessitates developing and establishing standardized FSL benchmarks in these domains.
6.7 Fully generalizable FSL